# Numerical instability of non-Hermitian Hamiltonian evolution

### v2

@@ I Introduction
Mathematically, EPs and NHSE arise from the unique properties of non-Hermitian matrices. EPs are special points in the parameter space where two or more eigenvalues and their corresponding eigenvectors coalesce, rendering the matrix defective [8]. Systems near EPs exhibit extreme sensitivity to parameter variations [1]. Beyond EPs, NHSE represents another key mechanism of extreme sensitivity. Two hallmark signatures of NHSE are the accumulation of a macroscopic number of eigenstates near the system’s boundary and a dramatic change in the energy spectrum under different boundary conditions [20, 21, 22, 23]. This phenomenon is closely linked to the spectral instability of non-Hermitian matrices. In general, if a non-Hermitian matrix $H$ is perturbed by an arbitrary matrix $\Delta$, the resulting spectral shift is not necessarily bounded by the norm $\lVert\Delta\rVert$ [1]. This instability poses significant challenges for numerical computations. Previous studies have shown that the spectrum obtained via numerical diagonalization can be unreliable due to extreme sensitivity to lattice size and numerical precision [24, 25]. To address this issue, the Generalized Brillouin Zone (GBZ) theory has been developed for analytically characterizing the spectrum of one-dimensional non-Hermitian systems in the thermodynamic limit [2, 20, 24].
@@
The rest of the paper is organized as follows. In Sec. II, we introduce the basic knowledge of the non-normal matrices, which theoretically demonstrates that strong non-normality can induce spectrum instability and unconventional dynamics. Next, we analytically solve the Hatano-Nelson and symplectic Hatano-Nelson models in Sec. III, whose condition number and pseudospectra are also obtained. Sec. IV presents the energy spectra and wave functions of the Hatano-Nelson and symplectic Hatano-Nelson model with different numerical precisions, highlighting the numerical instability during diagonalization. Sec. V provides the results of non-Hermitian evolution for both models at various numerical precisions, further revealing the instability of non-Hermitian evolution. In Sec. VI, we address a common misunderstanding, namely that a reliable spectrum alone does not guarantee the reliability of non-Hermitian evolution, while the accuracy of the wave functions is more critical. Finally, Sec. VII concludes the paper and offers several perspectives for future work. Appendix A introduces the Gaussian state simulation method and various observables of interest. In Appendix B, we monitor the evolution of the $U$ matrix to observe the manifestation of numerical errors directly. Appendix C provides additional numerical data to further support the findings presented in the main text.
@@ II Fundamentals of the non-normal matrix
In this work, we consider the case of a non-Hermitian Hamiltonian, where $H\neq H^{\dagger}$. Specifically, we primarily focus on non-normal matrices, which demand $[H,H^{\dagger}]\neq 0$. For any normal matrix, the spectral theorem guarantees the existence of a unitary transformation $U$ that diagonalizes the matrix, i.e., $H=U\Lambda U^{-1}$, where $\Lambda$ is a diagonal matrix representing the energy spectrum, and the columns of $U$ correspond to the eigenstates. However, if $H$ is non-normal but still diagonalizable, there exists an invertible (albeit non-unitary) matrix $V$ such that $H=V\Lambda V^{-1}$. Due to the non-unitarity of $V$, both left and right eigenvectors are required to construct a biorthogonal basis. In extreme cases, such as at exceptional points (EPs), the non-Hermitian Hamiltonian becomes non-diagonalizable, meaning its geometric multiplicity is lower than its algebraic multiplicity. In such cases, the Hamiltonian can only be transformed into the Jordan normal form.
@@
Besides the spectrum instability, the strong non-normality can induce counterintuitive dynamics [37]. Naively, considering the evolution of the $\lVert e^{-\mathrm{i}Ht}\rVert$, because $\lVert e^{-\mathrm{i}Ht}\rVert=\lVert Ve^{-\mathrm{i}\Lambda t}V^{-1}\rVert\leq\text{cond}(V)\cdot\lVert e^{-\mathrm{i}\Lambda t}\rVert$, if the $H$ is a normal matrix, i.e. $\text{cond}(V)=1$, the evolution of $\lVert e^{-\mathrm{i}Ht}\rVert$ is determined by the spectrum of the $H$ as expected. However, for a highly non-normal matrix of very large $\text{cond}(V)$, the transient dynamics can greatly deviate from the behavior of $\lVert e^{-\mathrm{i}\Lambda t}\rVert$. As we will demonstrate in Sec. V, for the Hatano-Nelson model, $\lVert e^{-\mathrm{i}Ht}\rVert$ initially grows exponentially with time before stabilizing at late times, as predicted by the $\lVert e^{-\mathrm{i}\Lambda t}\rVert$. Interestingly, we will show that this sharp early-time growth of $\lVert e^{-\mathrm{i}Ht}\rVert$ is closely tied to the numerical instability of the non-Hermitian evolution.
@@ III.2 Symplectic Hatano-Nelson model
Figure 4: The numerical diagonalized wave functions of the Hatano-Nelson with different system sizes. The figure includes all wave functions. $\gamma=0.8$. The blue line is the numerical result of double precision, while the red line is the numerical result of the digit 50. The dashed green line characterizes the theoretical exponential localization of the Hatano-Nelson model with localization length $\xi=1/\text{In}\sqrt{(J+\gamma)/(J-\gamma)}$. Its form is $|\psi(x)|=ce^{x/\xi}$, $c$ varies with system size. It should be noted that the green lines are not the exact results of wave functions; they only approximately capture the exponential localization behaviors of the wave functions. When $L\geq 40$, the majority of wave functions with double precision display a larger localization length than the theoretical value, while the localization lengths with digit 50 fit with the theoretical value. We also note that for double precision, some eigenstates still obey the theoretical localization length. However, owing to the restrictions of the double precision, once the absolute value of the wave function decreases to around $10^{-16}\sim 10^{-15}$, it will get smooth.
@@ IV Numerical errors of diagonalization
Interestingly, the diagonalization results for $\gamma=-0.8$ (see Fig. 3(b)) are significantly more numerically stable than those for $\gamma=0.8$ (see Fig. 3(a)), despite both non-Hermitian matrices being identical after a transpose operation. This discrepancy stems from the numerical diagonalization algorithm. To illustrate this, we consider the standard eigenvalue computation method: $QR$ algorithm. The $QR$ algorithm follows these steps: first, it performs the $QR$ decomposition $H=Q_{0}R_{0}$, and then multiplies the $QR$ factors in the reverse order to obtain $H_{1}=R_{0}Q_{0}$. The algorithm then iterates this factor-and-reverse process $H_{1}=Q_{1}R_{1}$, $H_{2}=R_{1}Q_{1}$, $H_{2}=Q_{2}R_{2}\dots$. This process eventually leads to an upper triangular matrix, whose diagonal entries are the eigenvalues of $H$. Therefore, this algorithm requires many times of $QR$ decompositions, which decomposes a matrix into the product of an orthogonal matrix and an upper triangular matrix. A preliminary observation is that for $\gamma=-0.8$, the upper triangular elements of the Hamiltonian $H_{\text{HN}}$ are larger than those in the lower triangular. In other words, the Hamiltonian of $\gamma=-0.8$ is closer to the upper triangular matrix than $\gamma=0.8$. Consequently, the computation for $\gamma=-0.8$ may converge much faster than $\gamma=0.8$, thus accumulating fewer errors during iterations. However, given the complexity of the diagonalization algorithm, this argument remains speculative. Furthermore, as shown in Fig. S2, we observe that the numerical results of wave functions for $\gamma=-0.8$ can be far more accurate than $10^{-16}$, even with double precision. The exact reason for this remarkable accuracy in the case of negative $\gamma$ remains elusive and warrants further investigation. We emphasize that for large system size $L$, such as $L=160$ in Fig. 3(b), the energy eigenvalues will also become complex, i.e., incorrect, which means the numerical instability still exists for $\gamma=-0.8$.
@@
which describes the exponential decay of wave function with localization length $\xi=1/\text{log}(r)=1/\text{log}\sqrt{(J+\gamma)/(J-\gamma)}$. Under the normalization condition $\||E^{(m)}_{\text{OBC}}\rangle\rVert^{2}=1$, which is automatically enforced by most numerical diagonalization algorithms, the absolute values of wave function components near the boundary can become much smaller than $10^{-16}$ for large $L$. This precision limit implies that eigenstates obtained with double precision become unreliable. Indeed, as shown in Fig. 4, when $L\gtrsim 40$, the numerically obtained wave functions deviate from theoretical predictions. These incorrect results can be categorized into two types: 1) Some wave functions decay exponentially with the expected rate $1/\xi$ until their amplitudes reach approximately $10^{-16}$, after which they remain at that level rather than continuing to decay (see also a similar pattern in spectral decomposition of variational quantum states [54]). 2) Other wave functions decay exponentially but at a rate smaller than the predicted $1/\xi$, ensuring that the smallest wave function amplitude at the left boundary remains close to $10^{-16}$, the approximate lower bound of double precision. When numerical precision is increased to 50-digit precision, this lower bound improves to at least $10^{-50}$, significantly smaller than the absolute values of the wave function at all sites. Consequently, as displayed in Fig. 4, the wave functions obtained with 50-digit precision roughly agree with analytical results. Moreover, we also improve the precision from double to digit $20$ and digit $30$. As shown in Fig. S4 and S5, the diagonalized wave functions get closer and closer to the analytical results. Interestingly, we will demonstrate in Sec. V that the incorrectness of the diagonalized wave functions is closely related to the incorrectness of the non-Hermitian evolution. Even though this incorrectness of the wave functions is extremely small, the incorrectness of the corresponding non-Hermitian evolution can be significant.
@@ V Numerical errors of non-Hermitian evolution
The two models introduced in section III are $U(1)$-symmetric and non-interacting, ensuring the preservation of Gaussianity. Their non-Hermitian evolution has been previously simulated using Gaussian-state methods [35]. In this study, we also employ the Gaussian-state simulation method from Ref. [35] for comparison. The observables of interest—correlation, density distribution, particle current, and entanglement entropy can be calculated using the Gaussian state simulation method. A brief overview of the simulation method and the definitions of these observables are provided in Appendix A.
@@
In terms of entanglement, because the numerically correct steady state tends to be more frozen, namely, the particles can scarcely move due to the Pauli exclusion principle. It is reasonable to predict that the correct entanglement should be smaller than the results presented in Ref. [35]. Indeed, as shown in Fig. 10, the entanglement entropy computed with 50-digit precision is consistently smaller than that obtained with double precision, except for $L=20$, where double precision is already sufficient for accurate results. Therefore, the qualitative conclusion from Ref. [35] that the steady-state entanglement obeys an area law for $\gamma\neq 0$ remains valid.
@@
Akin to the Hatano-Nelson model, the numerical results of the symplectic Hatano-Nelson model in Fig. 8, 9, 10, 11 of Ref. [35] (including density, correlation, current, entanglement) are also incorrect due to numerical errors. Firstly, as indicated in Fig. 11 and 12, the middle smooth varying region disappears when the numerical precision is improved from double to digit $50$. Moreover, as shown in Fig. 14, the current in chain $A$ or $B$ with digit 50 (i.e. the spin current in Fig. 10(b), 10(c), 10(d) of Ref. [35]) vanishes after long-time averaging, in contrast to the results with double precision, where a non-zero current exists in the late time. Furthermore, Fig. 14(b) shows that the steady-state entanglement at digit $50$ in the skin regime ($\gamma>\delta$) is smaller than the results with double precision as well. It is vital to note that we only present the results of $\gamma>\delta$ for the symplectic Hatano-Nelson model since the results of $\gamma<\delta$ are numerically stable owing to the disappearance of the skin effect, i.e., small condition number.
@@
In addition, we also calculate the $\lVert e^{-\mathrm{i}ht}U\rVert$, where $h$ and $U$ are the matrix representations of the Hamiltonian and initial state, respectively (see Appendix. A). We observe that its behavior closely resembles that of $\lVert e^{-\mathrm{i}Ht}\rVert$. This indicates that the non-Hermitian evolution $\dfrac{e^{-\mathrm{i}Ht}|\psi_{0}\rangle}{\lVert e^{-\mathrm{i}Ht}|\psi_{0}\rangle\rVert}$ can be numerically challenging. Specifically speaking, as time $t$ increases, $\lVert e^{-\mathrm{i}Ht}|\psi_{0}\rangle\rVert$ exponentially grows until it approximately approaches the value of $\text{cond}(V)$. For highly non-normal systems, $\text{cond}(V)$ can be extremely large, causing the elements of $\dfrac{e^{-\mathrm{i}Ht}|\psi_{0}\rangle}{\lVert e^{-\mathrm{i}Ht}|\psi_{0}\rangle\rVert}$ to become much smaller than the default double precision, leading to numerical errors. Furthermore, during the timescale in which $\lVert e^{-\mathrm{i}Ht}\psi_{0}\rVert$ is smaller than about $O(10^{16})$, we can predict that the non-Hermitian evolution is numerically correct with double precision. For example, as shown in Fig. 8 and 14, the current and the entanglement entropy with double precision agree with the results of digit $50$ for an initial period. In summary, the large condition number not only induces the spectrum instability as illustrated in Sec. IV but also contributes to the numerical instability of the non-Hermitian evolution.

### v3

@@ I Introduction
Mathematically, EPs and NHSE arise from the unique properties of non-Hermitian matrices. EPs are special points in the parameter space where two or more eigenvalues and their corresponding eigenvectors coalesce, rendering the matrix defective [8]. Systems near EPs exhibit extreme sensitivity to parameter variations [1]. Beyond EPs, NHSE represents another key mechanism of extreme sensitivity. Two hallmark signatures of NHSE are the accumulation of a macroscopic number of eigenstates near the system’s boundary and a dramatic change in the energy spectrum under different boundary conditions [20, 21, 22, 23]. This phenomenon is closely linked to the spectral instability of non-Hermitian matrices. In general, if a non-Hermitian matrix $H$ is perturbed by an arbitrary matrix $\Delta$, the resulting spectral shift is not necessarily bounded by the norm $\lVert\Delta\rVert$ [1, 24]. This instability poses significant challenges for numerical computations. Previous studies have shown that the spectrum obtained via numerical diagonalization can be unreliable due to extreme sensitivity to lattice size and numerical precision [25, 26]. To address this issue, the Generalized Brillouin Zone (GBZ) theory has been developed for analytically characterizing the spectrum of one-dimensional non-Hermitian systems in the thermodynamic limit [2, 20, 25].
@@
The rest of the paper is organized as follows. In Sec. II, we introduce the basic knowledge of the non-normal matrices, which theoretically demonstrates that strong non-normality can induce spectrum instability and unconventional dynamics. Next, we analytically solve the Hatano-Nelson and symplectic Hatano-Nelson models in Sec. III, whose condition number and pseudospectra are also obtained. Sec. IV presents the energy spectra and wave functions of the Hatano-Nelson and symplectic Hatano-Nelson model with different numerical precisions, highlighting the numerical instability during diagonalization. Sec. V provides the results of non-Hermitian evolution for both models at various numerical precisions, further revealing the instability of non-Hermitian evolution. In Sec. VI, we address a common misunderstanding, namely that a reliable spectrum alone does not guarantee the reliability of non-Hermitian evolution, while the accuracy of the wave functions is more critical. Finally, Sec. VII concludes the paper and offers several perspectives for future work. Appendix A provides evidence that the observed numerical asymmetry arises from the $QR$ algorithm. Appendix B introduces the Gaussian state simulation method and various observables of interest. In Appendix C, we monitor the evolution of the $U$ matrix to observe the manifestation of numerical errors directly. Appendix D provides additional numerical data to further support the findings presented in the main text. The generalizations to the disordered and interacting systems are briefly illustrated in Appendix E.
@@ II Fundamentals of the non-normal matrix
In this work, we consider the case of a non-Hermitian Hamiltonian, where $H\neq H^{\dagger}$. Specifically, we primarily focus on non-normal matrices, which demand $[H,H^{\dagger}]\neq 0$. For any normal matrix, the spectral theorem guarantees the existence of a unitary transformation $U$ that diagonalizes the matrix, i.e., $H=U\Lambda U^{-1}$, where $\Lambda$ is a diagonal matrix representing the energy spectrum, and the columns of $U$ correspond to the eigenstates. However, if $H$ is non-normal but still diagonalizable, there exists an invertible (albeit non-unitary) matrix $V$ such that $H=V\Lambda V^{-1}$. Due to the non-unitarity of $V$, both left and right eigenvectors are required to construct a biorthogonal basis. In extreme cases, such as at EPs, the non-Hermitian Hamiltonian becomes non-diagonalizable, meaning its geometric multiplicity is lower than its algebraic multiplicity. In such cases, the Hamiltonian can only be transformed into the Jordan normal form.
@@
Besides the spectrum instability, the strong non-normality can induce counterintuitive dynamics [24]. Naively, considering the evolution of the $\lVert e^{-\mathrm{i}Ht}\rVert$, because $\lVert e^{-\mathrm{i}Ht}\rVert=\lVert Ve^{-\mathrm{i}\Lambda t}V^{-1}\rVert\leq\text{cond}(V)\cdot\lVert e^{-\mathrm{i}\Lambda t}\rVert$, if $H$ is a normal matrix, i.e. $\text{cond}(V)=1$, the evolution of $\lVert e^{-\mathrm{i}Ht}\rVert$ is determined by the spectrum of $H$ as expected. However, for a highly non-normal matrix of very large $\text{cond}(V)$, the transient dynamics can greatly deviate from the behavior of $\lVert e^{-\mathrm{i}\Lambda t}\rVert$. As we will demonstrate in Sec. V, for the Hatano-Nelson model, $\lVert e^{-\mathrm{i}Ht}\rVert$ initially grows exponentially with time before stabilizing at late times, as predicted by the $\lVert e^{-\mathrm{i}\Lambda t}\rVert$. Interestingly, we will show that this sharp early-time growth of $\lVert e^{-\mathrm{i}Ht}\rVert$ is closely tied to the numerical instability of the non-Hermitian evolution.
In summary, as shown in Table 1, both spectral perturbations and transient dynamics can be significantly amplified in non-Hermitian systems exhibiting the NHSE, indicating the potential for numerical instability.
Table 1: Summary of four key concepts for non-normal matrices: non-normality, condition number, spectral instability, and dynamical instability.
Concept Key Expression
Non-normality $[H,H^{\dagger}]\neq 0$
Condition number $\text{cond}(V)=\lVert V\rVert\cdot\lVert V^{-1}\rVert=\dfrac{s_{\max}(V)}{s_{\min}(V)}$
Spectral instability $\{z\,|\,\mathrm{dist}(z,\sigma(H))<\varepsilon\cdot\text{cond}(V)\}$
Dynamical instability $\lVert e^{-\mathrm{i}Ht}\rVert\leq\text{cond}(V)\cdot\lVert e^{-\mathrm{i}\Lambda t}\rVert$
@@ III.2 Symplectic Hatano-Nelson model
Figure 4: The numerical diagonalized wave functions of the Hatano-Nelson with different system sizes. The figure includes all wave functions for $\gamma=0.8$. The blue lines are the numerical results of double precision, while the red lines are the numerical results of the digit 50. The dashed green lines characterize the theoretical exponential localization of the Hatano-Nelson model with localization length $\xi=1/\ln\sqrt{(J+\gamma)/(J-\gamma)}$. Its form is $|\psi(x)|=ce^{x/\xi}$, $c$ varies with system size. It should be noted that the green lines are not the exact results of wave functions; they only approximately capture the exponential localization behaviors of the wave functions. When $L\geq 40$, the majority of wave functions with double precision display a larger localization length than the theoretical value, while the localization lengths with digit 50 closely match the theoretical value. We also observe that for double precision, some eigenstates still obey the theoretical localization length. However, owing to the restrictions of the double precision, once the absolute value of the wave function decreases to around $10^{-16}\sim 10^{-15}$, it will get smooth.
@@ IV Numerical errors of diagonalization
Interestingly, the diagonalization results for $\gamma=-0.8$ (see Fig. 3(b)) are significantly more numerically stable than those for $\gamma=0.8$ (see Fig. 3(a)), despite both non-Hermitian matrices being identical after a transpose operation. This discrepancy stems from the numerical diagonalization algorithm. To illustrate this, we consider the standard eigenvalue computation method: the $QR$ algorithm. The $QR$ algorithm follows these steps: first, it performs the $QR$ decomposition $H=Q_{0}R_{0}$, and then multiplies the $QR$ factors in the reverse order to obtain $H_{1}=R_{0}Q_{0}$. The algorithm then iterates this factor-and-reverse process $H_{1}=Q_{1}R_{1}$, $H_{2}=R_{1}Q_{1}$, $H_{2}=Q_{2}R_{2}\dots$, which eventually leads to a quasi-upper-triangular matrix, composed of $1\times 1$ blocks on the diagonal for real eigenvalues and $2\times 2$ blocks for complex conjugate pairs of eigenvalues [54]. In Appendix A, we simulate the diagonalization routine and demonstrate that the numerical asymmetry originates from the $QR$ algorithm. Specifically, we find that the matrices $H_{1},H_{2},H_{3},\dots$ generated during the $QR$ iterations exhibit distinct structural differences between $\gamma=0.8$ and $\gamma=-0.8$. For $\gamma=0.8$, the $QR$ decomposition will produce extremely small denominators during the Householder reflections, leading to numerical instability. In contrast, for $\gamma=-0.8$, such small denominators are avoided, resulting in improved numerical stability. Furthermore, as shown in Fig. S2, we observe that the numerical results of wave functions for $\gamma=-0.8$ can be far more accurate than $10^{-16}$, even with double precision. The exact reason for this remarkable accuracy in the case of negative $\gamma$ remains elusive and warrants further investigation. We emphasize that for large system size $L$, such as $L=160$ in Fig. 3(b), the energy eigenvalues will also become complex, i.e., incorrect, which means the numerical instability still exists for $\gamma=-0.8$.
@@
which describes the exponential decay of wave function with localization length $\xi=1/\ln(r)=1/\ln\sqrt{(J+\gamma)/(J-\gamma)}$. Under the normalization condition $\||E^{(m)}_{\text{OBC}}\rangle\rVert^{2}=1$, which is automatically enforced by most numerical diagonalization algorithms, the absolute values of wave function components near the boundary can become much smaller than $10^{-16}$ for large $L$. This precision limit implies that eigenstates obtained with double precision become unreliable. Indeed, as shown in Fig. 4, when $L\gtrsim 40$, the numerically obtained wave functions deviate from theoretical predictions. These incorrect results can be categorized into two types: 1) Some wave functions decay exponentially with the expected rate $1/\xi$ until their amplitudes reach approximately $10^{-16}$, after which they remain at that level rather than continuing to decay (see also a similar pattern in spectral decomposition of variational quantum states [55]). 2) Other wave functions decay exponentially but at a rate smaller than the predicted $1/\xi$, ensuring that the smallest wave function amplitude at the left boundary remains close to $10^{-16}$, the approximate lower bound of double precision. When numerical precision is increased to 50-digit precision, this lower bound improves to at least $10^{-50}$, significantly smaller than the absolute values of the wave function at all sites. Consequently, as displayed in Fig. 4, the wave functions obtained with 50-digit precision roughly agree with analytical results. Moreover, we also improve the precision from double to digit $20$ and digit $30$. As shown in Fig. S4 and S5, the diagonalized wave functions get closer and closer to the analytical results. Interestingly, we will demonstrate in Sec. V that the incorrectness of the diagonalized wave functions is closely related to the incorrectness of the non-Hermitian evolution. Even though this incorrectness of the wave functions is extremely small, the incorrectness of the corresponding non-Hermitian evolution can be significant.
@@ V Numerical errors of non-Hermitian evolution
The two models introduced in section III are $U(1)$-symmetric and non-interacting, ensuring the preservation of Gaussianity. Their non-Hermitian evolution has been previously simulated using Gaussian-state methods [36]. In this study, we also employ the Gaussian-state simulation method from Ref. [36] for comparison. The observables of interest—correlation, density distribution, particle current, and entanglement entropy can be calculated using the Gaussian state simulation method. A brief overview of the simulation method and the definitions of these observables are provided in Appendix B.
@@
In terms of entanglement, because the numerically correct steady state tends to be more frozen—namely, the particles can scarcely move due to the Pauli exclusion principle—it is reasonable to predict that the correct entanglement should be smaller than the results presented in Ref. [36]. Indeed, as shown in Fig. 10, the entanglement entropy computed with 50-digit precision is consistently smaller than that obtained with double precision, except for $L=20$, where double precision is already sufficient for accurate results. Therefore, the qualitative conclusion from Ref. [36] that the steady-state entanglement obeys an area law for $\gamma\neq 0$ remains valid.
@@
Akin to the Hatano-Nelson model, the numerical results of the symplectic Hatano-Nelson model in Figs. 8, 9, 10, 11 of Ref. [36] (including density, correlation, current, entanglement) are also incorrect due to numerical errors. Firstly, as indicated in Fig. 11 and 12, the middle smooth varying region disappears when the numerical precision is improved from double to digit $50$. Moreover, as shown in Fig. 14, the current in chain $A$ or $B$ with digit 50 (i.e. the spin current in Fig. 10(b), 10(c), 10(d) of Ref. [36]) vanishes after long-time averaging, in contrast to the results with double precision, where a non-zero current exists in the late time. Furthermore, Fig. 14(b) shows that the steady-state entanglement at digit $50$ in the skin regime ($\gamma>\delta$) is smaller than the results with double precision as well. It is vital to note that we only present the results of $\gamma>\delta$ for the symplectic Hatano-Nelson model since the results of $\gamma<\delta$ are numerically stable owing to the disappearance of the skin effect, i.e., small condition number.
@@
In addition, we also calculate the $\lVert e^{-\mathrm{i}ht}U\rVert$, where $h$ and $U$ are the matrix representations of the Hamiltonian and initial state, respectively (see Appendix. B). We observe that its behavior closely resembles that of $\lVert e^{-\mathrm{i}Ht}\rVert$. This indicates that the non-Hermitian evolution $\dfrac{e^{-\mathrm{i}Ht}|\psi_{0}\rangle}{\lVert e^{-\mathrm{i}Ht}|\psi_{0}\rangle\rVert}$ can be numerically challenging. Specifically speaking, as time $t$ increases, $\lVert e^{-\mathrm{i}Ht}|\psi_{0}\rangle\rVert$ exponentially grows until it approximately approaches the value of $\text{cond}(V)$. For highly non-normal systems, $\text{cond}(V)$ can be extremely large, causing the elements of $\dfrac{e^{-\mathrm{i}Ht}|\psi_{0}\rangle}{\lVert e^{-\mathrm{i}Ht}|\psi_{0}\rangle\rVert}$ to become much smaller than the default double precision, leading to numerical errors. Furthermore, during the timescale in which $\lVert e^{-\mathrm{i}Ht}\psi_{0}\rVert$ is smaller than about $O(10^{16})$, we can predict that the non-Hermitian evolution is numerically correct with double precision. For example, as shown in Fig. 8 and 14, the current and the entanglement entropy with double precision agree with the results of digit $50$ for an initial period. In summary, the large condition number not only induces the spectrum instability as illustrated in Sec. IV but also contributes to the numerical instability of the non-Hermitian evolution.